OpenStudy (anonymous):

pleasee help!! stuck on this question! Decide on what substitution to use, and then evaluate the given integral using a substitution. x(x-6)^1/4 dx

OpenStudy (anonymous):

So far I did U=x-6 du= 1 dx so dx= 1du ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

your x wont cancel like you are probably expecting, so you'll need to solve for it in terms of u.

OpenStudy (anonymous):

so I got then 1 (integral symbol) u^1/4 +1 du =1 u^5/4/(5/4) +c =1 (x-6)^5/4/(5/4) + C

OpenStudy (anonymous):

where did I go wrong :/?

OpenStudy (anonymous):

\[\int\limits_{}^{}(u-6)(\sqrt{u})du\]

OpenStudy (anonymous):

where did the square root come from?

OpenStudy (anonymous):

im sorry I read the 1/4 as 1/2

OpenStudy (anonymous):

oh okay no worries. so its ∫(u−6)(u)du

OpenStudy (anonymous):

an I plug in the "u" ?

OpenStudy (anonymous):

yes with u^(1/4)

OpenStudy (anonymous):

so would it be (x-6)(x-6)^1/4 du ?

OpenStudy (anonymous):

then multiply the u's --> u^(5/4)-6u^(1/4)

OpenStudy (anonymous):

Then split the integral

OpenStudy (anonymous):

so -6u^3/2

OpenStudy (anonymous):